0=y^2-8x+4y-20 ive tried everything but i can't get the answer. its asking for the equation in standard form,vertex,focus,directrix,and the functions in terms of y.please help i need to know this to pass my math class.
Answer by stanbon(75887)

its asking for the equation in standard form, vertex, focus, directrix,and the functions in terms of y.
----------------
y^2-8x+4y-20 = 0
-------------
y^2+4y+? = 8x-20 + ?
------
y^2+4y+4 = 3x-20+4
(y+2)^2 = 3x-16
3x = (y+2)^2+16
x = (1/3)(y+2)^2 + (16/3)
------
vertex: (-2,16/3)
---
4p = 1/3
p = 1/12
directrix: x = -2-(1/12) ; x = -25/12
focus: -2+1/12 = 23/12 :::Ans: (23/12 , 16/3)
-------------
Cheers,
Stan H.
===============